OpenStudy (anonymous):

does the sum from n=0 to infinity of (7^n)/(9^(n+1)) converge or diverge

OpenStudy (sidsiddhartha):

it converges

OpenStudy (anonymous):

what did u do to sole that like what test

OpenStudy (anonymous):

i can tell it converges but by what test? integral seems as though it would be too complicated

OpenStudy (sidsiddhartha):

vn=1/3^(2n+2) it will produce a finite value

OpenStudy (anonymous):

can you explain that please

OpenStudy (irishboy123):

\[7^n/(9^(n+1)=\] \[7^n/(9^n).9=\] \[(1/9).(7/9)^n\] IOW the individual terms tend to zero meaning it could converge. next step integrate the series as if it were a function and if the sum < infinity, it does converge. Integrate {(1/)(7/9)^n} dn and you should get (1/9)(1/n)((7/9)^(n=1)), 7^2/ 9^3 or note that the ratio between successive terms is < 1 so it converges under the Ratio test... make sense....?